OpenStudy (anonymous):

For fun:- which values of n, this statement is true \(5|n^7-4\)

ganeshie8 (ganeshie8):

\[n^7 \equiv 4\pmod{5}\]

ganeshie8 (ganeshie8):

\[n^4\cdot n^3 \equiv 4\pmod{5}\] \[1\cdot n^3 \equiv 4\pmod{5}\] \[ n^3 \equiv 4\pmod{5}\]

OpenStudy (anonymous):

nice , keep going

ganeshie8 (ganeshie8):

take log both sides

ganeshie8 (ganeshie8):

\[\mathrm{ind}~ n^3 \equiv \mathrm{ind}~4 \pmod{5} \]

ganeshie8 (ganeshie8):

\[3~\mathrm{ind}~ n \equiv \mathrm{ind}~4 \pmod{5} \]

OpenStudy (anonymous):

what id "ind" :O

OpenStudy (mathmath333):

indian log

ganeshie8 (ganeshie8):

haha it is "log" equivalent for congruences

OpenStudy (anonymous):

lol not kidding !

ganeshie8 (ganeshie8):

let me finish this problem first maybe, after that il define what it is

OpenStudy (anonymous):

ok go ahead

ganeshie8 (ganeshie8):

\[3~\mathrm{ind}~ n \equiv \mathrm{ind}~4 \pmod{\phi(5)} \] \[3~\mathrm{ind}~ n \equiv 2 \pmod {4}\] \[\mathrm{ind}~ n \equiv 2 \pmod {4}\] \[ n \equiv 4 \pmod 5\]

ganeshie8 (ganeshie8):

so all the possible solutions are \(n = 4 + 5k\) where \(k\) is any integer

OpenStudy (anonymous):

nice ! ok now evaluate ind for me , cuz i dint got it

ganeshie8 (ganeshie8):

Alright, before talking about log/ind, lets do this without using this fancy stuff

ganeshie8 (ganeshie8):

\[n^3 \equiv 4\pmod{5}\] we can solve this using just algebra (division algorithm)

ganeshie8 (ganeshie8):

every integer can be represented in one of the forms : {5k, 5k+1, 5k+2, 5k+3, 5k+4} consider 5 cases and see which ones give you a remainder 0

OpenStudy (anonymous):

i used mod only to solve it xD n=0,1,2,3,4 mod 5 n^7=0,1,2,3,4 mod 5 n=4 mod 5 give the solution

ganeshie8 (ganeshie8):

ikr :) just wanted to show you these cool stuff

OpenStudy (anonymous):

yeah i wanna see these cool stuff :O

ganeshie8 (ganeshie8):

we can discuss after exams maybe..

ganeshie8 (ganeshie8):

ind is similar to log

OpenStudy (anonymous):

okkk

ganeshie8 (ganeshie8):

it follows all log properties : \[a^{\text{ind}~ n} \equiv n \] \[\text{ind} ~ n^k \equiv k~\text{ind} ~ n \] \[\text{ind} ~ (ab) \equiv ~\text{ind} ~ (a) +\text{ind} ~ (b) \] etc.. this is simple theory based on primitive roots. we can do more problems after exams :)

OpenStudy (anonymous):

:O

OpenStudy (mathmath333):

google cant find ind symbol

ganeshie8 (ganeshie8):

i can make you understand everything about indices and primitive roots in less than 1 hour mathmath

ganeshie8 (ganeshie8):

lets do it if you're interested and when you're free as i also want to understand these more

OpenStudy (mathmath333):

m free now by the way

ganeshie8 (ganeshie8):

Good, we will start with fermat's little theorem

ganeshie8 (ganeshie8):

Fermat's little theorem tells us \[a^{p-1}\equiv 1 \pmod{p}\] whenever \(a\) is not divisible by \(p\) Notice that here we have \(p-1\) numbers relatively prime to \(p\) : \(\{1,2,3,...,p-1\}\)

OpenStudy (mathmath333):

yes i knoww ,ok

ganeshie8 (ganeshie8):

each of these numbers satisfy the fermat's little theorem, for example : \[2^{4}\equiv 1 \pmod{5}\]

ganeshie8 (ganeshie8):

Also we have \[4^{4}\equiv 1 \pmod{5}\]

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{(2^{12}-1)~(\mod 13)\equiv 0\hspace{.33em}\\~\\ (2^{18}-1)~(\mod 19)\equiv 0\hspace{.33em}\\~\\ }\end{align}\)

ganeshie8 (ganeshie8):

im sure you knew all these, we're going to define primitive root based on this

OpenStudy (mathmath333):

yes i know feramet

ganeshie8 (ganeshie8):

look closely at \[4^{4}\equiv 1 \pmod{5}\]

ganeshie8 (ganeshie8):

look more closely at the exponent is \(4\) the least exponent for which aboveyields true ?

ganeshie8 (ganeshie8):

Nope. because we have \[4^{\color{Red}{2}} \equiv 1 \pmod {5}\]

OpenStudy (mathmath333):

yes it works for it works also for \(\large\tt \begin{align} \color{black}{(4^{2}(\mod 5)\equiv 1\hspace{.33em}\\~\\ }\end{align}\)

ganeshie8 (ganeshie8):

yes so the exponent "4" is not least there is another number "2" for which you also get 1, yes ?

OpenStudy (mathmath333):

yes

ganeshie8 (ganeshie8):

how about the powers of 2

ganeshie8 (ganeshie8):

\[2^4 \equiv 1 \pmod{5}\]

ganeshie8 (ganeshie8):

is there any power less than 4 that gives you 1 ?

OpenStudy (mathmath333):

no

OpenStudy (mathmath333):

\(2^{0}\) does gives

ganeshie8 (ganeshie8):

we are only considering the powers between 1 and p-1

ganeshie8 (ganeshie8):

powers between 1 and 4 here

OpenStudy (mathmath333):

not it doen'st give 2^1=2 (mod 5) 2^2=4 (mod 5) 2^3=3 (mod 5)

ganeshie8 (ganeshie8):

\(2^1 \equiv 2\) \(2^2 \equiv 4\) \(2^3 \equiv 8\equiv 3\) \(2^4 \equiv 16 \equiv 1 \) so the \(p-1\) is the least power that gives you 1 for 2

ganeshie8 (ganeshie8):

we say \(2\) is a primitive root of \(5\)

ganeshie8 (ganeshie8):

now that you have seen what a primitive root is, lets formally define it

OpenStudy (mathmath333):

2 is primitive root because below 2^4 doesnt gives \(\equiv 1\)

OpenStudy (mathmath333):

for mod 5

ganeshie8 (ganeshie8):

2 is primitive root because only 2^4 gives 1

ganeshie8 (ganeshie8):

4 is not primitive root because 4^2 also gives 1

ganeshie8 (ganeshie8):

we are talking about mod 5 ofcourse

OpenStudy (mathmath333):

ok i see the defn

ganeshie8 (ganeshie8):

Definition : \(a\) is a primitive root of \(\mod p\) if \(p-1\) is the least value that satisfies \(a^{p-1}\equiv 1 \pmod{p}\)

OpenStudy (mathmath333):

ok i get that

ganeshie8 (ganeshie8):

lets do 2-3 examples before moving on to indices

ganeshie8 (ganeshie8):

find all the primitive roots of \(\mod 5\)

OpenStudy (mathmath333):

how do i find that ?

OpenStudy (mathmath333):

2 is of cos

ganeshie8 (ganeshie8):

you're in mod 5, so you just need to check 5 numbers : {1, 2, 3, 4, 5}

OpenStudy (mathmath333):

3 should be

ganeshie8 (ganeshie8):

Right, 2 and 3 are primitive roots of 5 are there any more ?

OpenStudy (mathmath333):

no cuz 4 is a multiple of 2

ganeshie8 (ganeshie8):

do you mean, 4 is not a primitive root because \(4^2 \equiv 1 \pmod{5}\)

OpenStudy (mathmath333):

yes also \(4^2 \equiv 1 \pmod{5}\) is same as \(2^4 \equiv 1 \pmod{5}\)}

ganeshie8 (ganeshie8):

Ahh I see! thats a good observation

ganeshie8 (ganeshie8):

last example on primitive roots : find all the primitive roots of 7

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{2,3,5\hspace{.33em}\\~\\ }\end{align}\)

ganeshie8 (ganeshie8):

thats fast! how did u get them ?

OpenStudy (mathmath333):

all are prime below \(7\)

ganeshie8 (ganeshie8):

2 is not a primitive root of 7 because \(2^3 \equiv 1 \pmod{7}\)

ganeshie8 (ganeshie8):

you need to stick to the definition it doesn't matter whether a number is prime or not

OpenStudy (mathmath333):

ok so if \(a^n \equiv 1 mod~ p\) a is not primitive to p

ganeshie8 (ganeshie8):

\(a\) is not a primitive root only if that congruence is true for some \(n\) less than \(p-1\)

ganeshie8 (ganeshie8):

2 is not a primitive root because we have \(2^3 \equiv 1 \pmod{7}\) and 3 is less than 6

ganeshie8 (ganeshie8):

how ever 3 is a primitive root because \(3^6\equiv 1 \pmod{7}\) and there are no other powers that give you 1

ganeshie8 (ganeshie8):

it takes some time to get use to the definition

OpenStudy (mathmath333):

ok so \(2^4 \equiv 1 \pmod{5}\) 2 is primitive beacause \(4\not <5-1\) for \(mod ~5\)

ganeshie8 (ganeshie8):

2 is primitive root because there are no other powers that give you 1 4 is the least power

OpenStudy (mathmath333):

yes i see it will take time

ganeshie8 (ganeshie8):

here are the primitive roots for few prime mods : ``` 5 : {2, 3} 7 : {3, 5} 11 : {2, 6, 7, 8} 13 : {2, 6, 7, 11} ```

ganeshie8 (ganeshie8):

notice 11 and 13 have four primitive roots each

ganeshie8 (ganeshie8):

it is true that every prime has atleast one primitive root

OpenStudy (anonymous):

since its satisfies Fermat !

OpenStudy (mathmath333):

so there are primitive roots only for prime numbers ?

OpenStudy (anonymous):

yesss

ganeshie8 (ganeshie8):

thats a good question, the answer is no. lets move to indices for now

OpenStudy (mathmath333):

lol

OpenStudy (anonymous):

haha the answer is yes